If 200g of water at 20 degrees Celsius is combined with 100g of water at 80 degrees Celsius, what is the final temperature of the 200g of water?
1
Expert's answer
2011-02-07T08:18:40-0500
The final temperature of the 200g of water would be equal to the final temperature of the whole (200+100) mixture:
According to the heat balance equation,
Cm1Δ T1 = Cm2Δ T2
Δ T1 = (T-20)
Δ T2 = (80-T)
200*(T-20) = 100* (80 - T)
2T-40 = 80 - T
3T = 120
T = 40 C.